# TIGER: A Generating-Then-Ranking Framework for Practical Python Type Inference

### v1

@@ Abstract
Python’s dynamic typing system offers flexibility and expressiveness but can lead to type-related errors, prompting the need for automated type inference to enhance type hinting. While existing learning-based approaches show promising inference accuracy, they struggle with practical challenges in comprehensively handling various types, including complex generic types and (unseen) user-defined types.
In this paper, we introduce TIGER, a two-stage generating-then-ranking (GTR) framework, designed to effectively handle Python’s diverse type categories. TIGER leverages fine-tuned pre-trained code models to train a generative model with a span masking objective and a similarity model with a contrastive training objective. This approach allows TIGER to generate a wide range of type candidates, including complex generics in the generating stage, and accurately rank them with user-defined types in the ranking stage. Our evaluation on the ManyTypes4Py dataset shows TIGER’s advantage over existing methods in various type categories, notably improving accuracy in inferring user-defined and unseen types by 11.2% and 20.1% respectively in Top-5 Exact Match. Moreover, the experimental results not only demonstrate TIGER’s superior performance and efficiency, but also underscore the significance of its generating and ranking stages in enhancing automated type inference.
@@ I Introduction
However, these learning-based approaches face limitations in considering various visible types, notably generic types and (unseen) user-defined types. Note that user-defined types in this paper also include third-party types. Classification-based approaches [12, 11, 13] rely on fixed-sized type vocabularies (e.g., 1,000 types for TypeWriter [12]), leading to numerous out-of-vocabulary (OOV) types, such as complex generic types or unseen user-defined types. Similarity-based approaches [1, 2, 17] use import information to gather user-defined types and rank candidate types based on contextual similarities. However, these approaches struggle to cover the myriad generic types (e.g.,Union[str,int]”) formed by elementary types (e.g.,Union”, “str”, “int”). Generation-based approaches [14, 15, 16, 3] can generate creative types, including complex generics, but most struggle with recognizing and predicting unseen user-defined types not in the training set.
@@
To overcome these limitations, we introduce a two-stage framework named generating-then-ranking (GTR framework), as depicted in Figure 1. This framework combines the strengths of both generation-based and similarity-based approaches, allowing for the consideration of comprehensive visible types. In the initial generating stage, our objective is to produce diverse candidate types, including commonly used elementary types like “UUID” and generic types like “Union[str,int]”, to fill the type placeholders (i.e., <TYPE> ” in Figure 1). Subsequently, in the ranking stage, visible user-defined types, potentially unseen during training, such as “IDMap” and “IDMapKey”, are effectively ranked alongside the generated candidates based on similarities. Furthermore, for efficiency and cost considerations, we opt for relatively lightweight pre-trained code models like the 220M-parameter CodeT5 [14], specifically fine-tuned for type inference, diverging from the use of LLMs like ChatGPT.
@@ II-A Rule-based Type Inference
Rule-based methods in type inference rely on predefined rules for inferring variable types. A rule can only be activated when all the underlying premises are known, after which it determines the type of the variable based on a set conclusion [3]. To address the imperative for static type hints within dynamically typed programming languages, numerous techniques have emerged for the purpose of type inference and verification. Notable examples include Pyright and Pylance (Microsoft) [8], Pyre (Meta) [7], pytype (Google) [9], and Python’s official type checker, Mypy [6]. Apart from industry tools, several academic methodologies have emerged for type inference across various programming languages like Python and JavaScript [28, 29, 30, 31, 32]. While these approaches excel in terms of accuracy, they confront a significant challenge stemming from dynamic language features and external function calls [17]. This challenge manifests as a limitation in their coverage, hindering their ability to provide comprehensive type inference scenarios.
@@ II-B Learning-based Type Inference
Similarity-based Approaches. Allamanis et al. [1] presented a graph neural network model that leverages probabilistic reasoning and deep similarity learning to predict types, including rare and user-defined types, based on a program’s structure, names, and patterns. Mir et al. [2] proposed Type4Py, a hierarchical neural network model employing deep similarity learning to infer likely types for program elements by distinguishing between similar and dissimilar types in a high-dimensional space. Peng et al. introduced HiTyper, a hybrid type inference approach combining static inference and similarity-based models by leveraging type dependency graphs (TDGs) to record and integrate type dependencies among variables, facilitating iterative static inference and neural predictions until the complete inference of TDGs. While these approaches demonstrate the capability to handle arbitrary types, they require pre-determined candidates, limiting their effectiveness in adequately considering numerous generic types.
@@
Generation-based Approaches. Following the type annotation conventions recommended in PEPs, the Python type inference problem can be transformed into a conditional generation problem. For instance, it can be treated as a cloze-style fill-in-blank task where a type placeholder is inserted after a variable that requires annotation. Code pre-trained models [21, 14, 15, 16] are then employed to generate types fill in the placeholder with types based on their learned code naturalness. To fully leverage their generation capability, some studies [34] have fine-tuned these pre-trained models to improve inference accuracy. These generation-based approaches, with fill-in-blank task formulation, lack awareness of unseen user or library-defined types. In a recent study, Peng et al. [3] introduced a generation-based approach known as TypeGen for Python type inference, which relies on parameter-forzen LLMs like gpt-3.5 and gpt-4. TypeGen integrates type dependencies derived from lightweight static analysis with in-context learning [35] to construct few-shot Chain-of-Thought (CoT) prompts. These prompts are employed to generate both type predictions and accompanying explanations. While this LLM-based generative approach demonstrates notable effectiveness in type inference, it faces practical challenges, particularly with regard to the scalability concerns arising from the resource-intensive interactions with expensive LLMs.
II-C Code Pre-trained Models in SE
@@
Recent research [36, 14, 37] indicates that code pre-trained models encapsulate valuable information about code syntax, identifier and namespace concepts, and natural language naming. To harness this rich code information for downstream software engineering tasks, researchers commonly fine-tune these models on task-specific data. Examples of such tasks include code search [38], AST tagging [36], and vulnerability repairing [39]. Researchers have explored the application of prompt-based approaches with code pre-trained models in various software engineering tasks. For instance, Wang et al. [40] investigated the effectiveness of prompt learning in code intelligence tasks, such as clone detection and code summarization. Another study by Wang et al. [41] utilized code pre-trained models as knowledge bases and employed prompt learning for variable explanation. In this study, we strategically fine-tune code pre-trained models to facilitate practical type inference.
@@ III Problem Definition
We define the type inference problem within a Python function as a cloze-style fill-in-blank task [21, 42, 14, 16], focusing on three categories of variables: local variables, function arguments, and return types. To annotate the type for a target variable, we insert a type placeholder based on its category, adhering to the annotation guidelines outlined in the PEPs. In Figure 2, the three numbered “ <TYPE> ” represent the type placeholders inserted for local variables, function arguments, and return types, respectively. In practical applications, we address one target variable in a given Python function at a time, following the approach adopted in previous studies [12, 2, 3]. Functions containing type placeholders are referred to as type-missed functions, where suitable types need to be predicted to fill the placeholders with difference categories of types, including builtin types, generic types, and use-defined or third-party types.
@@ IV Approach
The overview of TIGER is presented in Figure 3. TIGER first offline trains a generation model and a similarity model by fine-tuning existing code pre-trained models (i.e., base models) using annotated Python functions (i.e., ① and ②). The trained models are used to support the online two-stage generating-then-ranking inference (i.e., ③).
@@ IV-C2 Contrastive Fine-tuning
  • Negative Instance Construction. Constructing suitable negative instances for the expected type is crucial for effective contrastive training. TIGER achieves this by generating $K$ negative instances from two distinct sources. For the given type-missed function (i.e., anchor), the first source utilizes the trained generation model, employing a beam search algorithm during model decoding to generate $K$ candidates (cf. Section IV-B3). These candidates typically encompass builtin types, commonly used third-party types, and generic types. The second source involves user-defined types obtained through import analysis. TIGER combines the types from both sources, excluding the positive instance, and randomly selects $K$ ones as the negative instances.
@@ IV-D2 Type Ranking
TIGER conducts import analysis, as detailed in Section IV-C1, to acquire user-defined types available for the type placeholder in $func$. These user-defined types, combined with the generated candidates, constitute a more comprehensive candidate pool. Note that if a generated candidate is neither a built-in type nor a generic type whose base is a built-in type, and it is not found in the user or library-defined types, we exclude it from the candidate pool. Subsequently, TIGER ranks all candidates in the pool by considering both generative likelihood and contextual similarity for each candidate. For a given candidate $cand$ in the pool, TIGER computes its generative likelihood using the generation model (cf. Section IV-B3) and its contextual similarity with $func$ using the similarity model (cf. Section IV-C3).
@@ V Evaluation
  • RQ1 (Effectiveness): How effectively can TIGER infer types for variables? How effectively can TIGER handle with various categories of types, especially generic types and (unseen) user defined types?
@@
  • RQ3 (Efficiency): How effectively can TIGER infer types for different sources of types, encompassing elementary types, generic types, and (unseen) user-defined types?
@@ V-A2 Dataset
In line with prior studies [2, 17], we conduct the evaluation on the ManyTypes4Py dataset [27], which is partitioned into training and test sets with an 8:2 project ratio (no overlapping projects between training set and test set), resulting in 242,954 instances for training and 85,205 instances for testing. For the TypeGen evaluation, the authors further sampled 10,000 instances from the entire test set to reduce the computational cost associated with calling large language models (LLMs). To maintain consistency in our experimental setup, we employ this sampled subset [53] as the definitive test set for all experiments pertaining to RQ1-RQ4. Among the test instances involving user-defined types, 579 instances require the inference of types that are previously unseen in training set. These unseen types are challenging to predict and provide insights into the practicality of the inference approach. The training set is utilized for training both TIGER and baselines, as well as providing few-shot examples for the prompt construction in TypeGen. The test set is then employed to assess the type inference effectiveness of TIGER and baselines. Table I provides an overview of the training and test sets, illustrating the distributions of different type and variable categories. “Ele”, “Gen”, and “Usr (Unseen)” represent different type categories [3], denoting Elementary, Generic, and (unseen) user-defined Types, respectively. “Var”, “Arg”, and “Ret” represent different variable categories [3], denoting Local Variables, Function Arguments, and Return Values, respectively.
@@ V-A3 Baselines
Ele Gen Usr (Unseen) All Ele Gen Usr (Unseen) All Ele Gen Usr (Unseen) All
@@ V-B1 Overall Results
Across the entire test set (Column “All”), our approach, TIGER, demonstrates Top-1, 3, and 5 Exact Match metrics of 86.2%, 93.2%, and 94.6%, respectively, and Top-1, 3, and 5 Base Match metrics of 92.3%, 96.2%, and 97.4%, respectively. Considering fine-grained type categories (Columns “Ele”, “Gen”, and “Usr (Unseen)”), TIGER achieves either the best or second-best results in terms of Exact Match and Base Match. Particularly noteworthy are the substantial improvements (see Rows “$\Delta$ compared with the best baseline”) in almost all type categories, especially in unseen user-defined types (up by 20.1% and 23.4% in Top-5 Exact and Base Match, respectively), with minimal declines (less than 1%) observed in few type categories such as Top-1 Exact Match for elementary types (Ele).
An interesting observation is that the Base Match metrics of all the approaches are notably higher than the Exact Match metrics for generic types (i.e., Column “Gen”) compared with the other two type categories Columns “Ele” and “Usr (Unseen)”). This discrepancy arises because inference approaches correctly predict the base type “Base” but face challenges predicting the specific combinations of “[T1,T2,...]”.
@@ V-B2 Comparison with Classification-based Approach
The superiority is attributed to the limitations of classification-based approaches, such as TypeWriter, which rely on fixed-sized type vocabularies. Such vocabularies often fail to comprehensively cover diverse generic types and user-defined types in their visible candidate set. TIGER addresses this limitation by leveraging the capacity of its generating stage to produce complex generic types and the ranking capability of its ranking stage to handle (unseen) user-defined types. Notably, while TypeWriter achieves about 30%-40% accuracy for unseen types, as its 1,000-type vocabulary covers a portion of unseen types in the ManyType4Py test set, it is unable to predict types beyond its vocabulary limit.
@@ V-B3 Comparison with Similarity-based Approaches
While demonstrating comparable accuracy for elementary types, TIGER significantly outperforms the similarity-based baseline approaches, Type4Py and HiTyper, for generic types and (unseen) user-defined types in terms of Exact Match and Base Match metrics. Specifically, in Top-5 Exact Match, TIGER achieves notable improvements over Type4Py and HiTyper, with enhancements of 17.8% and 18.8% respectively across the entire test set, ranging from 21.6% to 679.8% over Type4Py and from 31.4% to 47.9% over HiTyper for generic types and user-defined types.
The enhancements achieved by TIGER can be attributed to the following reasons. While similarity-based approaches can handle commonly used generic types (e.g.,list[str]”), they struggle with inferring diverse generic types due to the vast number of possible combinations. In contrast, TIGER leverages the creative type generation capacity of its generation model during the generating stage to overcome this limitation. Moreover, in the case of (unseen) user-defined types, TIGER’s superior performance is primarily attributed to its robustness and generalizability in the ranking stage, facilitated by the contrastive training objective [23].

### v3

@@ Abstract
Python’s dynamic typing system offers flexibility and expressiveness but can lead to type-related errors, prompting the need for automated type inference to enhance type hinting. While existing learning-based approaches show promising inference accuracy, they struggle with practical challenges in comprehensively handling various types, including complex parameterized types and (unseen) user-defined types.
In this paper, we introduce TIGER, a two-stage generating-then-ranking (GTR) framework, designed to effectively handle Python’s diverse type categories. TIGER leverages fine-tuned pre-trained code models to train a generative model with a span masking objective and a similarity model with a contrastive training objective. This approach allows TIGER to generate a wide range of type candidates, including complex parameterized types in the generating stage, and accurately rank them with user-defined types in the ranking stage. Our evaluation on the ManyTypes4Py dataset shows TIGER’s advantage over existing methods in various type categories, notably improving accuracy in inferring user-defined and unseen types by 11.2% and 20.1% respectively in Top-5 Exact Match. Moreover, the experimental results not only demonstrate TIGER’s superior performance and efficiency, but also underscore the significance of its generating and ranking stages in enhancing automated type inference.
Index Terms:
type inference, pre-trained code models, generating-then-ranking, contrastive learning
@@ I Introduction
However, these learning-based approaches face limitations in considering various visible types, notably parameterized types and (unseen) user-defined types. Note that user-defined types in this paper also include third-party types. Classification-based approaches [12, 11, 13] rely on fixed-sized type vocabularies (e.g., 1,000 types for TypeWriter [12]), leading to numerous out-of-vocabulary (OOV) types, such as complex parameterized types (e.g.,Dict[str,List[int]]”) or unseen user-defined types. Similarity-based approaches [1, 2, 17] use import information to gather user-defined types and rank candidate types based on contextual similarities. However, these approaches struggle to cover the myriad parameterized types (e.g.,Union[str,int]”) formed by elementary types (e.g.,Union”, “str”, “int”). Generation-based approaches [14, 15, 16, 3] can generate creative types, including complex parameterized types, but most struggle with recognizing and predicting unseen user-defined types not in the training set.
@@
To overcome these limitations, we introduce a two-stage framework named generating-then-ranking (GTR framework), as depicted in Figure 1. This framework combines the strengths of both generation-based and similarity-based approaches, allowing for the consideration of comprehensive visible types. In the initial generating stage, our objective is to produce diverse candidate types, including commonly used elementary types like “UUID” and parameterized types like “Union[str,int]”, to fill the type placeholders (i.e., <TYPE> ” in Figure 1). Subsequently, in the ranking stage, visible user-defined types, potentially unseen during training, such as “IDMap” and “IDMapKey”, are effectively ranked alongside the generated candidates based on similarities. Furthermore, for efficiency and cost considerations, we opt for relatively lightweight pre-trained code models like the 220M-parameter CodeT5 [14], specifically fine-tuned for type inference, diverging from the use of LLMs like ChatGPT.
@@ II-A Rule-based Type Inference
Rule-based methods in type inference rely on predefined rules for inferring variable types. A rule can only be activated when all the underlying premises are known, after which it determines the type of the variable based on a set conclusion [3]. To address the imperative for static type hints within dynamically typed programming languages, numerous techniques have emerged for the purpose of type inference and verification. Notable examples include Pyright and Pylance (Microsoft) [8], Pyre (Meta) [7], pytype (Google) [9], and Python’s official type checker, Mypy [6]. Apart from industry tools, several academic methodologies have emerged for type inference across various programming languages like Python and JavaScript [28, 29, 30, 31, 32, 33]. While these approaches excel in terms of accuracy, they confront a significant challenge stemming from dynamic language features and external function calls [17]. This challenge manifests as a limitation in their coverage, hindering their ability to provide comprehensive type inference scenarios.
@@ II-B Learning-based Type Inference
Similarity-based Approaches. Allamanis et al. [1] presented a graph neural network model that leverages probabilistic reasoning and deep similarity learning to predict types, including rare and user-defined types, based on a program’s structure, names, and patterns. Mir et al. [2] proposed Type4Py, a hierarchical neural network model employing deep similarity learning to infer likely types for program elements by distinguishing between similar and dissimilar types in a high-dimensional space. Peng et al. introduced HiTyper, a hybrid type inference approach combining static inference and similarity-based models by leveraging type dependency graphs (TDGs) to record and integrate type dependencies among variables, facilitating iterative static inference and neural predictions until the complete inference of TDGs. While these approaches demonstrate the capability to handle arbitrary types, they require pre-determined candidates, limiting their effectiveness in adequately considering numerous parameterized types.
@@
Generation-based Approaches. Following the type annotation conventions recommended in PEPs, the Python type inference problem can be transformed into a conditional generation problem. For instance, it can be treated as a cloze-style fill-in-blank task where a type placeholder is inserted after a variable that requires annotation. Pre-trained code models [21, 14, 15, 16] are then employed to generate types fill in the placeholder with types based on their learned code naturalness. To fully leverage their generation capability, some studies [35] have fine-tuned these pre-trained models to improve inference accuracy. These generation-based approaches, with fill-in-blank task formulation, lack awareness of unseen user or library-defined types. In a recent study, Peng et al. [3] introduced a generation-based approach known as TypeGen for Python type inference, which relies on parameter-forzen LLMs like gpt-3.5 and gpt-4. TypeGen integrates type dependencies derived from lightweight static analysis with in-context learning [36] to construct few-shot Chain-of-Thought (CoT) prompts. These prompts are employed to generate both type predictions and accompanying explanations. While this LLM-based generative approach demonstrates notable effectiveness in type inference, it faces practical challenges, particularly with regard to the scalability concerns arising from the resource-intensive interactions with expensive LLMs.
II-C Pre-trained Code Models in SE
@@
Recent research [37, 14, 38] indicates that pre-trained code models encapsulate valuable information about code syntax and semantics, identifier and namespace concepts, and natural language naming. To harness this rich code information for downstream software engineering tasks, researchers commonly fine-tune these models on task-specific data. Examples of such tasks include code generation/completion [39, 40, 41, 42], code search [43, 44], defect detection [45], vulnerability detection/location [46], and program repair [47, 48]. Researchers have explored the application of prompt-based approaches with pre-trained code models in various software engineering tasks. For instance, Wang et al. [49] investigated the effectiveness of prompt learning in code intelligence tasks, such as clone detection and code summarization. Another study by Wang et al. [50] utilized pre-trained code models as knowledge bases and employed prompt learning for variable explanation. In this study, we strategically fine-tune pre-trained code models to facilitate practical type inference.
@@ III Problem Definition
We define the type inference problem within a Python function as a cloze-style fill-in-blank task [21, 51, 14, 16], focusing on three categories of variables: local variables, function arguments, and return types. To annotate the type for a target variable, we insert a type placeholder based on its category, adhering to the annotation guidelines outlined in the PEPs. In Figure 2, the three numbered “ <TYPE> ” represent the type placeholders inserted for local variables, function arguments, and return types, respectively. In practical applications, we address one target variable in a given Python function at a time, following the approach adopted in previous studies [12, 2, 3]. Functions containing type placeholders are referred to as type-missed functions, where suitable types need to be predicted to fill the placeholders with difference categories of types, including builtin types, parameterized types, and use-defined or third-party types.
@@ IV Approach
The overview of TIGER is presented in Figure 3. TIGER first offline trains a generation model and a similarity model by fine-tuning existing pre-trained code models (i.e., base models) using annotated Python functions (i.e., ① and ②). The trained models are used to support the online two-stage generating-then-ranking inference (i.e., ③).
@@ IV-C2 Contrastive Fine-tuning
  • Negative Instance Construction. Constructing suitable negative instances for the expected type is crucial for effective contrastive training. TIGER achieves this by generating $K$ negative instances from two distinct sources. For the given type-missed function (i.e., anchor), the first source utilizes the trained generation model, employing a beam search algorithm during model decoding to generate $K$ candidates (cf. Section IV-B3). These candidates typically encompass builtin types, commonly used third-party types, and parameterized types. The second source involves user-defined types obtained through import analysis. TIGER combines the types from both sources, excluding the positive instance, and randomly selects $K$ ones as the negative instances.
@@ IV-D2 Type Ranking
TIGER conducts import analysis, as detailed in Section IV-C1, to acquire user-defined types available for the type placeholder in $func$. These user-defined types, combined with the generated candidates, constitute a more comprehensive candidate pool. Note that if a generated candidate is neither a built-in type nor a parameterized type whose base is a built-in type, and it is not found in the user or library-defined types, we exclude it from the candidate pool. Subsequently, TIGER ranks all candidates in the pool by considering both generative likelihood and contextual similarity for each candidate. For a given candidate $cand$ in the pool, TIGER computes its generative likelihood using the generation model (cf. Section IV-B3) and its contextual similarity with $func$ using the similarity model (cf. Section IV-C3).
@@ V Evaluation
  • RQ1 (Effectiveness): How effectively can TIGER infer types for variables? How effectively can TIGER handle with various categories of types, especially parameterized types and (unseen) user defined types?
@@
  • RQ3 (Efficiency): How effectively can TIGER infer types for different sources of types, encompassing elementary types, parameterized types, and (unseen) user-defined types?
@@ V-A2 Dataset
In line with prior studies [2, 17], we conduct the evaluation on the ManyTypes4Py dataset [27], which is partitioned into training and test sets with an 8:2 project ratio (no overlapping projects between training set and test set), resulting in 242,954 instances for training and 85,205 instances for testing. For the TypeGen evaluation, the authors further sampled 10,000 instances from the entire test set to reduce the computational cost associated with calling large language models (LLMs). To maintain consistency in our experimental setup, we employ this sampled subset [62] as the definitive test set for all experiments pertaining to RQ1-RQ4. Among the test instances involving user-defined types, 579 instances require the inference of types that are previously unseen in training set. These unseen types are challenging to predict and provide insights into the practicality of the inference approach. The training set is utilized for training both TIGER and baselines, as well as providing few-shot examples for the prompt construction in TypeGen. The test set is then employed to assess the type inference effectiveness of TIGER and baselines. Table I provides an overview of the training and test sets, illustrating the distributions of different type and variable categories. “Ele”, “Par”, and “Usr (Unseen)” represent different type categories [3], denoting Elementary, Parameterized, and (unseen) user-defined Types, respectively. “Var”, “Arg”, and “Ret” represent different variable categories [3], denoting Local Variables, Function Arguments, and Return Values, respectively.
@@ V-A3 Baselines
Ele Par Usr (Unseen) All Ele Par Usr (Unseen) All Ele Par Usr (Unseen) All
@@ V-B1 Overall Results
Across the entire test set (Column “All”), our approach, TIGER, demonstrates Top-1, 3, and 5 Exact Match metrics of 86.2%, 93.2%, and 94.6%, respectively, and Top-1, 3, and 5 Base Match metrics of 92.3%, 96.2%, and 97.4%, respectively. Considering fine-grained type categories (Columns “Ele”, “Par”, and “Usr (Unseen)”), TIGER achieves either the best or second-best results in terms of Exact Match and Base Match. Particularly noteworthy are the substantial improvements (see Rows “$\Delta$ compared with the best baseline”) in almost all type categories, especially in unseen user-defined types (up by 20.1% and 23.4% in Top-5 Exact and Base Match, respectively), with minimal declines (less than 1%) observed in few type categories such as Top-1 Exact Match for elementary types (Ele).
An interesting observation is that the Base Match metrics of all the approaches are notably higher than the Exact Match metrics for parameterized types (i.e., Column “Par”) compared with the other two type categories Columns “Ele” and “Usr (Unseen)”). This discrepancy arises because inference approaches correctly predict the base type “Base” but face challenges predicting the specific combinations of “[T1,T2,...]”.
@@ V-B2 Comparison with Classification-based Approach
The superiority is attributed to the limitations of classification-based approaches, such as TypeWriter, which rely on fixed-sized type vocabularies. Such vocabularies often fail to comprehensively cover diverse parameterized types and user-defined types in their visible candidate set. TIGER addresses this limitation by leveraging the capacity of its generating stage to produce complex parameterized types and the ranking capability of its ranking stage to handle (unseen) user-defined types. Notably, while TypeWriter achieves about 30%-40% accuracy for unseen types, as its 1,000-type vocabulary covers a portion of unseen types in the ManyType4Py test set, it is unable to predict types beyond its vocabulary limit.
@@ V-B3 Comparison with Similarity-based Approaches
While demonstrating comparable accuracy for elementary types, TIGER significantly outperforms the similarity-based baseline approaches, Type4Py and HiTyper, for parameterized types and (unseen) user-defined types in terms of Exact Match and Base Match metrics. Specifically, in Top-5 Exact Match, TIGER achieves notable improvements over Type4Py and HiTyper, with enhancements of 17.8% and 18.8% respectively across the entire test set, ranging from 21.6% to 679.8% over Type4Py and from 31.4% to 47.9% over HiTyper for parameterized types and user-defined types.
The enhancements achieved by TIGER can be attributed to the following reasons. While similarity-based approaches can handle commonly used parameterized types (e.g.,list[str]”), they struggle with inferring diverse parameterized types due to the vast number of possible combinations. In contrast, TIGER leverages the creative type generation capacity of its generation model during the generating stage to overcome this limitation. Moreover, in the case of (unseen) user-defined types, TIGER’s superior performance is primarily attributed to its robustness and generalizability in the ranking stage, facilitated by the contrastive training objective [23].